# Different types of syntactic agreement recruit the same units within large language models

### v1

@@ 2 Related work
More recent work has moved from distributed representations to identifying sparse, functionally meaningful components. Lakretz et al. (2019, 2021) used targeted ablations in LSTMs to identify model units that are responsible for number agreement. Complementing these findings, Mueller et al. (2022) applied counterfactual interventions to neuron activations to study how models encode agreement dependencies, and Boguraev et al. (2025) used causal interventions to characterize shared mechanisms across filler-gap constructions. Marks et al. (2024) and Lepori et al. (2023) isolated minimal subgraphs within models—so-called sparse feature circuits—that support subject-verb agreement. Using a feature-based approach, Durrani et al. (2020) identified subsets of units predictive of performance on various syntactic tasks, such as part-of-speech tagging and chunking a text into constituents. Duan et al. (2025) introduced a syntactic selectivity index (SSI), which quantifies how strongly a model unit differentiates grammatical versus ungrammatical sentences for a given syntactic phenomenon relative to other phenomena based on activation magnitude. They then tracked how these high-SSI units emerge over the course of training, and showed that ablating high-SSI units increases perplexity on grammatical sentences.
Prior work on multilingual models has shown that different languages rely on shared representations of syntactic structure. At the level of output probabilities, Michaelov et al. (2023) and Arnett et al. (2025) have studied cross-lingual generalization using structural priming approaches. At the level of full model representations, (morpho)syntactic features are encoded in shared geometries and patterns of unit activations (Chi et al., 2020; Papadimitriou et al., 2021; Brinkmann et al., 2025), and at a more granular level, syntax-sensitive neurons and circuits generalize across languages (de Varda and Marelli, 2023; Stanczak et al., 2022; Ferrando and Costa-jussà, 2024; cf. Tang et al., 2024).
@@ 4 Results
To situate these findings, we evaluated our overlap scores against two controls. First, we quantified chance-level overlap by computing the expected 2-fold intersection if 1% of units were sampled uniformly at random in each fold—yielding an expected overlap of 0.01% of all LLM units (Random in Figure 1). Second, we extracted only the grammatical sentences from each phenomenon (1,000 sentences), re-assigned condition labels (so 500 of them become “ungrammatical”) to form 500 pseudo-minimal pairs, and separated the resulting pairs into two folds of 250 pairs each. We then averaged the cross-validated unit overlap across the 67 pseudo-phenomena. If the isolated units responded to non-syntactic regularities (e.g., BLiMP template effects) or were outliers with unusually high activation magnitude or variance (He et al., 2024), pseudo-overlap would be high. This control (BLiMP-Control) also shows very low overlap (1.03%), indicating that the consistency of identified syntactic units cannot be explained by trivial factors.
@@ 4.2 Do different syntactic phenomena recruit the same model units?
The results are shown in Figure 4: across most phenomena, the within-category overlaps are reliably higher than the cross-category overlaps ($t=2.07$, $p=0.0312$ via a two-sample Welch $t$-test). Interestingly, this pattern appears to be driven by the three agreement categories: determiner-noun agreement, subject-verb agreement, and anaphor agreement (average within-category overlaps: 68.91%, 40.29%, and 23.10%, respectively). Other categories (e.g., ellipsis and argument structure) show much smaller differences between the within-category vs. cross-category overlaps. And some categories (e.g., quantifiers and irregular forms) show minimal within-category overlap (in the presence of high within-phenomenon overlap; Figure 1), which suggests that these phenomena do not form meaningful categories within LLMs and instead rely on phenomenon-specific units. See Appendix A and B for results at 0.5% and 5% localization thresholds, which show the same patterns.
@@ 4.3 Examination of cross-phenomenon relationships in languages other than English
Building on the English findings, we next asked whether our main results concerning the cross-phenomenon relationships generalize to other languages. To do so, we used two additional benchmarks: RuBLiMP for Russian (an Indo-European language; Taktasheva et al., 2024) and SLING for Chinese (a Sino-Tibetan language; Song et al., 2022). Both benchmarks are conceptually and structurally similar to BLiMP: RuBLiMP contains 45 phenomena of 1,000 minimal pairs each, grouped into 12 theoretically-defined syntactic categories, while SLING contains 38 phenomena with 1,000 pairs grouped into 9 categories. We conducted the analyses with the Gemma model because it supports both languages.
@@ 5 Discussion
We have shown that LLMs contain units that are consistently recruited for individual syntactic phenomena, and that different types of syntactic agreement draw on overlapping subsets of those units, indicating that agreement constitutes a meaningful category within LLMs (as independently demonstrated in English, Chinese, and Russian). These agreement-selective units also show non-trivial overlap across languages, pointing to a shared pool of units that support agreement-related computations. Moreover, the number of shared units increases with syntactic similarity: agreement-related units overlap more between English and Russian than between English and Chinese—a finding we extended in a broader analysis of 57 languages.
These results have several implications. First, they rule out the possibility that LLMs contain generic syntax-processing units, i.e., units that support syntactic computations across diverse phenomena (cf. Boguraev et al., 2025). Instead, distinct model units represent different aspects of grammatical knowledge, even for phenomena that have been previously grouped on theoretical grounds (Chomsky, 1981). However, a striking exception to these phenomenon-idiosyncratic units is agreement-related phenomena. In English, Russian, and Chinese, the most coherent categories—those with high within-category and low cross-category overlaps—are agreement categories. Moreover, different agreement categories overlap with each other, suggesting a shared functional substrate for syntactic agreement in LLMs. Another important implication has to do with cross-linguistic unit overlap in multilingual models. The fact that more syntactically similar languages exhibit greater overlap suggests that multilingual models organize syntactic representations in a way that reflects cross-linguistic similarity. Future studies may attempt to replicate this finding in humans by examining syntactic processing in bilingual and multilingual speakers to see whether syntactic processing is more similar between similar languages.
Finally, these findings have implications for cognitive science and neuroscience more broadly: the patterns of cross-phenomenon overlap in LLMs motivate investigations of whether humans, for example, show greater neural overlap for some phenomena than others. Although many differences exist between human brains and LLMs, LLMs offer a powerful system for deriving hypotheses about the neural organization of linguistic knowledge (e.g. Jain et al., 2024; Tuckute et al., 2024).
@@
Limitations of our work include the reliance on top-nn% unit localization, which may overlook more distributed syntactic representations. Our analyses are conducted in the models’ native activation basis; alternate representations might reveal structure that is not captured here (e.g., in non-basis-aligned subspaces; Wu et al., 2023; Geiger et al., 2024; see Mueller et al., 2024 for an overview). Additionally, our analyses are primarily correlational: although we do demonstrate the causal involvement of the target units in Section §4.1 and Appendix G, further causal interventions and circuit-level analyses will be needed to evaluate the importance of shared agreement units to model performance on agreement phenomena. Another limitation is that we have focused on benchmarks that depend on comparing grammatical sentences to their ungrammatical counterparts. However, grammatical knowledge and processing go beyond detecting syntactic violations.
In the future, we plan to apply a similar approach to other syntactic phenomena—including sentences that use infrequent structures, contain ambiguities, or long-distance dependencies—to develop a more comprehensive understanding of how grammatical knowledge and processing are instantiated in LLMs. Moreover, future work could extend these experiments to other languages for which BLiMP-like minimal pair benchmarks have recently been made available, such as Dutch (BLiMP-NL; Suijkerbuijk et al., 2025) or Turkish (TurBLiMP; Başar et al., 2025). Another rich avenue for future work is understanding the relationship between syntax and semantics in model representations. We took an initial step toward this goal with BLiMP-Lex, demonstrating that the localized agreement units respond more reliably to syntactic than semantic violations, but disentangling the two remains challenging (Huang et al., 2021; Reddy and Wehbe, 2021; Zhang et al., 2022).

### v2

@@ 2 Related work
More recent work has moved from distributed representations to identifying sparse, functionally meaningful components. Lakretz et al. (2019, 2021) used targeted ablations in LSTMs to identify model units that are responsible for number agreement. Complementing these findings, Mueller et al. (2022) applied counterfactual interventions to neuron activations to study how models encode agreement dependencies, Boguraev et al. (2025) used causal interventions to characterize shared mechanisms across filler-gap constructions, and Lu et al. (2025) used interventions to improve factual recall accuracy in non-English languages. Marks et al. (2024) and Lepori et al. (2023) isolated minimal subgraphs within models—so-called sparse feature circuits—that support subject-verb agreement. Using a feature-based approach, Durrani et al. (2020) identified subsets of units predictive of performance on various syntactic tasks, such as part-of-speech tagging and chunking a text into constituents. Duan et al. (2025) introduced a syntactic selectivity index (SSI), which quantifies how strongly a model unit differentiates grammatical versus ungrammatical sentences for a given syntactic phenomenon relative to other phenomena based on activation magnitude. They then tracked how these high-SSI units emerge over the course of training, and showed that ablating high-SSI units increases perplexity on grammatical sentences.
Prior work on multilingual models has shown that different languages rely on shared representations of syntactic structure. At the level of output probabilities, Michaelov et al. (2023) and Arnett et al. (2025) have studied cross-lingual generalization using structural priming approaches. At the level of full model representations, (morpho)syntactic features are encoded in shared geometries and patterns of unit activations (Chi et al., 2020; Papadimitriou et al., 2021; Brinkmann et al., 2025), and at a more granular level, syntax-sensitive neurons and circuits generalize across languages (de Varda and Marelli, 2023; Stanczak et al., 2022; Ferrando and Costa-jussà, 2024; Zhang et al., 2025; cf. Tang et al., 2024).
@@ 4 Results
To situate these findings, we evaluated our overlap scores against two controls. First, we quantified chance-level overlap by computing the expected 2-fold intersection if 1% of units were sampled uniformly at random in each fold—yielding an expected overlap of 0.01% of all LLM units (Random in Figure 1). Second, we extracted only the grammatical sentences from each phenomenon (1,000 sentences), re-assigned condition labels (so 500 of them became “ungrammatical”) to form 500 pseudo-minimal pairs, and separated the resulting pairs into two folds of 250 pairs each. We then averaged the cross-validated unit overlap across the 67 pseudo-phenomena. If the isolated units responded to non-syntactic regularities (e.g., BLiMP template effects) or were outliers with unusually high activation magnitude or variance (He et al., 2024), pseudo-overlap would be high. This control (BLiMP-Control) also shows very low overlap (1.03%), indicating that the consistency of identified syntactic units cannot be explained by trivial factors.
@@ 4.2 Do different syntactic phenomena recruit the same model units?
The results are shown in Figure 4: across most phenomena, the within-category overlaps are reliably higher than the cross-category overlaps ($t=2.07$, $p=0.0312$ via a two-sample Welch $t$-test). Interestingly, this pattern appears to be driven by the three agreement categories: determiner-noun agreement, subject-verb agreement, and anaphor agreement (average within-category overlaps: 68.91%, 40.29%, and 23.10%, respectively). Other categories (e.g., ellipsis and argument structure) show much smaller differences between the within-category vs. cross-category overlaps. And some categories (e.g., quantifiers and irregular forms) show minimal within-category overlap (in the presence of high within-phenomenon overlap; Figure 1), which suggests that these phenomena do not form meaningful categories within LLMs and instead rely on phenomenon-specific units. See Appendixes A and B for results at 0.5% and 5% localization thresholds, which show the same patterns.
@@ 4.3 Examination of cross-phenomenon relationships in languages other than English
Building on the English findings, we next asked whether our main results concerning the cross-phenomenon relationships generalize to other languages. To do so, we used two additional benchmarks: RuBLiMP for Russian (an Indo-European language; Taktasheva et al., 2024) and SLING for Chinese (a Sino-Tibetan language; Song et al., 2022). Both benchmarks are conceptually and structurally similar to BLiMP: RuBLiMP contains 45 phenomena with 1,000 minimal pairs each, grouped into 12 theoretically-defined syntactic categories, while SLING contains 38 phenomena with 1,000 pairs grouped into 9 categories. We conducted the analyses with the Gemma model because it supports both languages.
@@ 5 Discussion
We have shown that LLMs contain units that are consistently recruited for individual syntactic phenomena, and that different types of syntactic agreement draw on overlapping subsets of those units, indicating that agreement constitutes a meaningful category within LLMs (as independently demonstrated in English, Chinese, and Russian). These agreement-selective units also show non-trivial overlap across languages, pointing to a shared pool of units that support agreement-related computations. Moreover, the number of shared units increases with syntactic similarity: agreement-related units overlap more between English and Russian than between English and Chinese (a finding we extended to 57 languages).
These results have several implications. First, they argue against the idea that LLMs contain generic syntax-processing units, i.e., units that support syntactic computations across diverse phenomena (cf. Boguraev et al., 2025). Instead, distinct model units represent different aspects of grammatical knowledge, even for phenomena that have been previously grouped on theoretical grounds (Chomsky, 1981). However, a striking exception to these phenomenon-idiosyncratic units is agreement-related phenomena. In English, Russian, and Chinese, the most coherent categories—those with high within-category and low cross-category overlaps—are agreement categories. Moreover, different agreement categories overlap with each other, suggesting a shared functional substrate for syntactic agreement in LLMs. Another implication has to do with cross-linguistic unit overlap in multilingual models. The fact that more syntactically similar languages exhibit greater overlap suggests that multilingual models organize syntactic representations in a way that reflects cross-linguistic similarity.
Finally, these findings have implications for cognitive science and neuroscience: the patterns of cross-phenomenon overlap in LLMs motivate investigations of whether humans, for example, show greater neural overlap for some phenomena than others, and whether bilingual or multilingual speakers exhibit more similar syntactic processing for more closely related languages. Although many differences exist between human brains and LLMs, LLMs offer a powerful system for deriving hypotheses about the neural organization of linguistic knowledge (e.g. Jain et al., 2024; Tuckute et al., 2024).
@@
Limitations of our work include the reliance on top-nn% unit localization, which may overlook more distributed syntactic representations. Our analyses are conducted in the models’ native activation basis; alternate representations might reveal structure that is not captured here (e.g., in non-basis-aligned subspaces; Wu et al., 2023; Geiger et al., 2024; see Mueller et al., 2024 for an overview). Additionally, our analyses are primarily correlational: although we do demonstrate the causal involvement of the target units in Section 4.1 and Appendix G, further causal interventions and circuit-level analyses will be needed to evaluate the importance of shared agreement units to model performance on agreement phenomena. Another limitation is that we have focused on benchmarks that depend on comparing grammatical sentences to their ungrammatical counterparts. However, grammatical knowledge and processing go beyond detecting syntactic violations.
In the future, we plan to apply a similar approach to other syntactic phenomena—including sentences that use infrequent structures, contain ambiguities, or long-distance dependencies—to develop a more comprehensive understanding of how grammatical knowledge and processing are instantiated in LLMs. Moreover, future work could extend these experiments to other languages for which BLiMP-like minimal pair benchmarks have recently been made available, such as Dutch (BLiMP-NL; Suijkerbuijk et al., 2025) or Turkish (TurBLiMP; başar-2025). Another rich avenue for future work is understanding the relationship between syntax and semantics in model representations. We took an initial step toward this goal with BLiMP-Lex, demonstrating that the localized agreement units respond more reliably to syntactic than semantic violations, but disentangling the two remains challenging (Huang et al., 2021; Reddy and Wehbe, 2021; Zhang et al., 2022).